OpenStudy (anonymous):

write the following equation in its equivalent logarithmic form 14^3=x

OpenStudy (unklerhaukus):

\[14^3=x\] taking the log of both sides \[\ln(14^3)=\ln x\] \[3\ln(14)=\ln x\]

OpenStudy (kinggeorge):

Alternatively,\[\log_{14} (x)=3\]is also correct if you take log base 14.

Directrix (directrix):

Logarithms are exponents. Attached is a definition of logs. Learn it so that you can easily move between the two forms.

Directrix (directrix):

In the expression 14^3=x, Three is the exponent. So, ______________ = 3 The base of the exponential expression 14^3 is 14. So, log base 14 ____ = 3. The only place left for the X is in the remaining blank. So, log base 14 of x = 3. The 14 is a subscript as written above by KingGeorge.